OpenStudy (anonymous):

when the equation a to the second power+10a+12=0 is solved using the quadratic formula one of the root is

OpenStudy (gw2011):

a^2+10a+12 x=[-b+/-sq rt (b^2-4ac)]2a a=1 b=10 c=12 x=[-10+/-sq rt (10^2-4(1)(12)]/2(1) x=[-10+/-sq rt (100-48)]/2 x=[-10+/-sq rt (52)]/2 x=[-10+/-sq rt (4)(13)]/2 x=[-10+/-2 (sq rt 13)]/2 x=-5+/-sq rt 13 x=-5+sq rt 13 and -5-sq rt 13